OpenStudy (anonymous):

Given that f(x)=2x^2 -3 and g(x) =x-1/2, find f o g

OpenStudy (saifoo.khan):

f o g = f(g(x)) = gx into fx. 2x^2 - 3 2(x-1/2)^2 - 3

OpenStudy (anonymous):

\[f(g(x))=2(x-1/2)^{2}-3\] \[=2\left( x ^{2}-x+1/4 \right)-3\] \[=2x ^{2}-2x+1/2-3\] \[=2x ^{2}-2x-5/2\] @Saifoo.khan, Am i correct?

OpenStudy (saifoo.khan):

Yes totally @rudi19